# Efficient 3D Bayesian Full Waveform Inversion and Analysis of Prior Hypotheses

### v1

@@ Abstract
Spatially 3-dimensional seismic full waveform inversion (3D FWI) is a highly nonlinear and computationally demanding inverse problem that constructs 3D subsurface seismic velocity structures using seismic waveform data. To characterise non-uniqueness in the solutions we demonstrate Bayesian 3D FWI using an efficient method called physically structured variational inference applied to 3D acoustic Bayesian FWI. The results provide reasonable posterior uncertainty estimates, at a computational cost that is only an order of magnitude greater than that of standard, deterministic FWI. Furthermore, we deploy variational prior replacement to calculate Bayesian solutions corresponding to different classes of prior information at low additional cost, and analyse those prior hypotheses by constructing Bayesian L-curves. This reveals the sensitivity of the inversion process to different prior assumptions. Thus we show that fully probabilistic 3D FWI can be performed at a cost that may be practical in small FWI problems, and can be used to test different prior hypotheses.
@@ 1 Introduction
Seismic full waveform inversion (FWI) uses both phase and amplitude information from observed waveform data to estimate compatible subsurface seismic velocity maps [7]. Since FWI is a high dimensional, strongly nonlinear and ill-posed inverse problem, finding the true Earth model and its corresponding uncertainties using standard inversion methods is currently impossible [2, 17, 10].
Monte Carlo sampling methods are commonly used to solve nonlinear Bayesian inverse problems and quantify uncertainties. However, applying conventional Monte Carlo methods to Bayesian FWI is computationally expensive due to the curse of dimensionality [3]. Variational inference uses optimisation rather than random sampling to solve Bayesian problems, and can be more efficient and easier to scale to high dimensions [16]. Recently, variational methods have been applied to reduce the computational cost of both Bayesian 2D [9, 13] and 3D FWI [11, 6], to reduce the computational cost.
We demonstrate Bayesian 3D FWI with substantially improved efficiency compared to previous methods by using physically structured variational inference [PSVI: 14]. This method has proven to provide relatively accurate inversion results including uncertainty estimates at greatly reduced computational cost in 2D FWI problems. To analyse, test and potentially select between different prior hypotheses, we use variational prior replacement [VPR: 12] to obtain inversion results corresponding to different prior pdf’s at negligible additional computational cost.
In what follows we first summarize methods, then present 3D FWI results. We compare inversion results obtained using different prior hypotheses by building Bayesian L-curves, then discuss implications of this work and draw conclusions.
@@ 2.1 Variational Bayesian Inversion
$p(\mathbf{m}|\mathbf{d}_{obs})=\dfrac{p(\mathbf{d}_{obs}|\mathbf{m})p(\mathbf{m})}{p(\mathbf{d}_{obs})}$ (1)
@@
In variational inference, an optimal variational distribution $q(\mathbf{m})$ is selected from a family of distributions to best approximate the posterior distribution $p(\mathbf{m}|\mathbf{d}_{obs})$. This is often accomplished by minimising the Kullback-Leibler (KL) divergence (a measure of difference) between the posterior and variational distributions [5], or equivalently maximising the evidence lower bound (ELBO) of $\log p(\mathbf{d}_{obs})$ defined as:
$\text{ELBO}[q(\mathbf{m})]=\mathbb{E}_{q(\mathbf{m})}[\log p(\mathbf{m},\mathbf{d}_{obs})-\log q(\mathbf{m})]$ (2)
This is an optimisation problem with a fully probabilistic result.
@@ 2.2 Physically Structured Variational Inference (PSVI)
PSVI is an efficient variational method that employs a transformed Gaussian distribution with a specific covariance structure to best fit the posterior distribution [14]. A Gaussian distribution $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ is characterised by a mean vector $\boldsymbol{\mu}$, and a covariance matrix $\boldsymbol{\Sigma}$ which is often re-parametrised using Cholesky factorisation $\boldsymbol{\Sigma}=\mathbf{L}\mathbf{L^{T}}$, where $\mathbf{L}$ is a lower triangular matrix.
PSVI captures only specific (important) posterior correlations in parameters $\mathbf{m}$, guided by prior knowledge of the physics controlling typical similar inverse problems. Specifically, our implementation of PSVI includes posterior correlations between pairs of locations that are spatially proximal, typically within one dominant wavelength in FWI problems. All other parameter correlations are ignored. This is accomplished by solving for specific off-diagonal elements of $\mathbf{L}$ during inversion, and fixing all other elements to be zero. This sparse structure for $\mathbf{L}$ and hence $\boldsymbol{\Sigma}$, requires a manageable number of hyperparameters to be optimised while still capturing what are typically the most significant correlations.
@@
Considering the computational cost of 3D Bayesian FWI, a top priority is to improve its efficiency. PSVI does so at the cost of some loss of generality in the solution found due to the Gaussian foundations for the posterior distribution. [14] demonstrated that the method nevertheless produces reasonable statistical information about the full, nonlinear posterior distribution.
@@ 2.3 Variational Prior Replacement
Say we have two different prior pdf’s $p_{old}(\mathbf{m})$ and $p_{new}(\mathbf{m})$ (hereafter subscripts old and new are used to denote the order in which the pdf’s are used). According to Bayes’ rule (equation 1), the two posterior probability distributions can be calculated by:
@@
In equation 3, we assume that the Bayesian problem given $\mathbf{d}_{obs}$ and prior pdf $p_{old}(\mathbf{m})$ is solved using variational inference giving an approximate posterior $q_{old}(\mathbf{m})$. The third and forth lines in equation 4 are obtained by substituting equation 3 into the second line. Term $p_{new}(\mathbf{m})/p_{old}(\mathbf{m})$ updates the solution that includes the old prior to a solution that contains the new one, and we set constant $k=p_{old}(\mathbf{d}_{obs})/p_{new}(\mathbf{d}_{obs})$. This implies that the new posterior distribution can be calculated from the old one by updating prior information post inversion by a simple multiplication [8].
Evaluating $k$ requires the two evidence terms to be calculated which is computationally intractable. Zhao and Curtis [12] estimated a second variational distribution $q_{new}(\mathbf{m})$ to approximate $p_{new}(\mathbf{m}|\mathbf{d}_{obs})$ (fifth line in equation 4) by minimising the KL divergence between the new posterior and $q_{new}(\mathbf{m})$, without requiring the value of $k$ to be known explicitly. This step is called variational prior replacement (VPR), and we use PSVI to perform VPR because its probability value is cheap and easy to evaluate for any model parameter values post inversion. Using PSVI to find $q_{old}(\mathbf{m})$ thus allows us to perform VPR efficiently to find $q_{new}(\mathbf{m})$ in equation 4 for alternative prior distributions.
@@
We apply PSVI to a synthetic 3D acoustic FWI problem. The true velocity model is a part of the 3D overthrust model, which contains $63\times 101\times 101$ grid cells in Z (depth), Y and X directions, respectively, with a cell size of 50 m in each direction (Figure 1a). We model 81 impulsive surface sources (red stars) and 10,201 receivers (white dots) with spacings of 500 m and 50 m, respectively: details in Appendix A.
We define the same Uniform prior distribution as Zhang et al. [11], and define an uncorrelated diagonal Gaussian likelihood function for variational Bayesian inversion (details in Appendix A.). The inversion is performed using 6 s long waveform data in bands referred to as the low (1 – 6 Hz), intermediate (1 – 10 Hz), and high frequency band (1 – 13 Hz). Optimizations for posterior pdf’s using data in intermediate and high frequency bands used results from the previous band as a starting point.
3.1 Inversion Results

### v2

# Efficient Bayesian Full Waveform Inversion and Analysis of Prior Hypotheses in 3D

@@ Abstract
Spatially 3-dimensional seismic full waveform inversion (3D FWI) is a highly nonlinear and computationally demanding inverse problem that constructs 3D subsurface seismic velocity structures using seismic waveform data. To characterise non-uniqueness in the solutions, we demonstrate Bayesian 3D FWI using an efficient method called physically structured variational inference, and apply it to 3D acoustic Bayesian FWI. The results provide reasonable posterior uncertainty estimates, at a computational cost that is only an order of magnitude greater than that of standard, deterministic FWI. Furthermore, we deploy variational prior replacement to calculate Bayesian solutions corresponding to different classes of prior information at low additional cost. The results obtained using prior information that models should be smooth show loop-like high uncertainty structures that are consistent with fully nonlinear inversion results presented previously. These structures disappear when smoothing is not imposed, so we conclude that they may be caused by smoothness constraints in tomographic problems. We further analyse a variety of prior hypotheses by constructing Bayesian L-curves, which reveal the sensitivity of the inversion process to different prior assumptions. To our knowledge, this is the first study that allows such prior hypotheses to be compared in probabilistic 3D FWI at feasible computational cost. This work shows that fully probabilistic 3D FWI can be performed and can be used to test different prior hypotheses, at a cost that may be practical, at least in small problems.
@@ 1 Introduction
Seismic full waveform inversion (FWI) uses both phase and amplitude information from observed waveform data to estimate compatible subsurface seismic velocity maps [40]. FWI is often implemented using deterministic methods through a gradient-based optimisation in which a measure of misfit between the observed and simulated seismograms is minimised iteratively. Given the highly nonlinear nature of the FWI problem, achieving a good initial model is crucial to avoid convergence to a locally optimal solution, and due to the ill-posed nature of most FWI problems, additional regularisation is also required to stabilise every inversion [59, 4, 1]. While this approach provides a single point estimate of a solution, unfortunately it can not readily be extended to estimate accurate uncertainty in the solutions, and so can not be used to assess risks in post imaging decision-making processes robustly [3, 12, 57, 48, 35].
Bayesian inference solves inverse problems within a probabilistic framework, in which a family of all potential solutions and their uncertainties are described by the so-called posterior probability distribution function (pdf). The posterior pdf is calculated using Bayes’ rule to update prior knowledge about model parameters with new information from observed data. Monte Carlo sampling methods are commonly used to sample the Bayesian posterior pdf, in order to quantify uncertainties. However, applying conventional Monte Carlo methods to Bayesian FWI is computationally expensive due to the many model parameters involved, which incurs the curse of dimensionality [11, 33, 34]. Efficient sampling techniques have been developed in attempts to address this issue, by either reducing the dimensionality (number of unknown parameters) of the sampling problem [30, 39, 20, 29] or using data-model gradient information to enhance the sampling efficiency [17, 58, 6, 60], or a combination of both [7]. Nevertheless, direct posterior sampling methods such as these remain a computational challenge.
Monte Carlo methods are expensive in part because no structure is imposed a priori on the sampling distribution; sampling algorithms must therefore be sufficiently flexible to adapt to, and explore, any pdf topography to any level detail. Variational inference offers an alternative to Monte Carlo methods to solve Bayesian problems. They search for an estimate of the posterior pdf that has a particular structure imposed a priori. This usually comes in the form of a family of pdf’s which is expected to contain an acceptable approximation to the posterior pdf. The best estimate of the posterior pdf within that family can then be found by optimisation rather than random sampling. Because the range of possible pdf estimates is infinitely smaller than the range of general pdf’s explored by Monte Carlo methods, variational methods can be more efficient and easier to scale to high dimensional problems. Recently, variational methods have been applied to 2D Bayesian FWI, and have been shown to reduce its computational cost while still providing acceptable estimates of the posterior distribution [47, 5, 43, 21, 52, 45, 37, 44, 49, 54].
This work concerns 3D FWI, which typically includes amongst the most computationally challenging problems in subsurface science. Nevertheless, Zhang et al. [50] applied three variational methods to a synthetic acoustic 3D Bayesian FWI problem, demonstrating the feasibility of finding solutions. Lomas et al. [26] and Walker et al. [41] then applied one of the methods to a field dataset collected with an airgun source and ocean bottom nodes over a salt body in the Gulf of Mexico. While there was no independent test of the solution quality, they estimated that uncertainties were higher around the boundaries of the salt body, as is generally expected in true seismic imaging solutions [15]. Hoffmann et al. [19] also performed 3D FWI and local uncertainty estimation using a source subsampling strategy and ensemble Kalman filter [14]. In most of these 3D FWI studies a spatial smoothing operator is employed a priori to improve the convergence rate of the inversion. However, the obtained posterior distribution may then underestimate uncertainties, which in part exclude the true solution [50]. Moreover, the type and strength of regularisation applied is effectively a subjective hypothesis, and is difficult to choose in practice because almost any form of regularity in the true Earth varies spatially, and so in principle each such hypothesis should be analysed and tested. Considering the huge computational cost of 3D FWI, analysing different hypotheses by comparing Bayesian inversion results that follow from each one is usually impractical.
We demonstrate efficient Bayesian 3D FWI and an approach to analyse different prior hypotheses. The novel contributions of this work can be summarised as follows:
  1. 1.
    We show Bayesian 3D FWI with substantially improved efficiency compared to previous methods. The improvement is attributed to the use of a recently introduced variational method: physically structured variational inference [PSVI: 52], which has proven to provide relatively accurate inversion results including uncertainty estimates at greatly reduced computational cost in 2D FWI problems. Our experiment demonstrates that with PSVI, fully nolinear 3D Bayesian FWI can be solved using only an order of magnitude more computation than traditional, deterministic FWI. To further improve the accuracy of the 3D FWI results reported in previous studies, we use a multiscale inversion strategy [10] by performing Bayesian FWI in three frequency bands using relatively low, intermediate, then high frequency data in an attempt to reduce cycle skipping.
  2. 2.
    To analyse, test and potentially select between different prior hypotheses, we apply a variational prior replacement (VPR) methodology [53] to obtain inversion results corresponding to different prior pdf’s at negligible additional computational cost. We compare VPR results with those obtained from independent Bayesian inversion, and find that VPR can not provide perfect estimates of posterior uncertainties due to the extreme complexity of this 3D FWI problem. We therefore introduce a method to fine tune the VPR results with a small amount of extra computation, which improves the results significantly.
  3. 3.
    We use the posterior pdf’s obtained using different prior distributions to construct a so-called Bayesian L-curve, from which we analyse different prior assumptions, and select one optimal choice among different hypotheses.
In the next section we summarise the methods deployed in this study. Next, we apply PSVI to 3D FWI and compare the inversion results with those from other variational and Monte Carlo methods, from which we identify two potential problems that affect the inversion accuracy. We then show how these issues can be addressed and present improved inversion results. In addition, we compare inversion results obtained using different prior hypotheses by building a Bayesian L-curve. Finally, we discuss implications of this work and draw conclusions.
@@ 2.1 Variational Bayesian Inversion
$p(\mathbf{m}|\mathbf{d}_{obs})=\dfrac{p(\mathbf{d}_{obs}|\mathbf{m})p(\mathbf{m})}{p(\mathbf{d}_{obs})},$ (1)
@@
In this work, we focus on variational inference, in which an optimal variational distribution $q(\mathbf{m})$ is selected from a family of predefined probability distributions, to best approximate the true but unknown posterior distribution $p(\mathbf{m}|\mathbf{d}_{obs})$. This is often accomplished by minimising the Kullback-Leibler (KL) divergence (a measure of difference) between the posterior and variational distributions [23], or equivalently maximising the evidence lower bound (ELBO) of $\log p(\mathbf{d}_{obs})$ defined as [8]:
$\text{ELBO}[q(\mathbf{m})]=\mathbb{E}_{q(\mathbf{m})}[\log p(\mathbf{m},\mathbf{d}_{obs})-\log q(\mathbf{m})],$ (2)
[…]
@@ 2.2 Physically Structured Variational Inference (PSVI)
PSVI is an efficient variational method that, in one implementation, employs a transformed Gaussian distribution with a specific covariance structure to best fit the posterior distribution [52]. A Gaussian distribution $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ is characterised by a mean vector $\boldsymbol{\mu}$, and a covariance matrix $\boldsymbol{\Sigma}$. To ensure positive semi-definiteness, the covariance matrix is often re-parametrised using a Cholesky factorisation $\boldsymbol{\Sigma}=\mathbf{L}\mathbf{L^{T}}$, where $\mathbf{L}$ is a lower triangular matrix. Variational distribution $q(\mathbf{m})$ can then be obtained by applying an invertible transform $f$ to the Gaussian distribution:
$\log q(\mathbf{m})=\log\mathcal{N}(\boldsymbol{\mu},\mathbf{L}\mathbf{L^{T}})-\log\left|\det[\partial_{\mathbf{m}}f^{-1}(\mathbf{m})]\right|$ (3)
Term $\det[\cdot]$ calculates the absolute value of the determinant of the Jacobian matrix $\partial_{\mathbf{m}}f^{-1}(\mathbf{m})$, which accounts for the volume change effected by $f$ [31]. This transform converts a Gaussian random variable defined in an unbounded space into the space of physical model parameters $\mathbf{m}$ which is often bounded [22].
Modelling a full covariance matrix requires $n(n+1)/2$ hyperparameters to construct $\mathbf{L}$ ($n$ being the number of model parameters - or the dimensionality of an inverse problem) which is infeasible in 3D FWI problems in which $n$ is typically of order greater than $10^{5}$. On the other hand, modelling only the diagonal elements (the commonly applied, so-called ‘mean-field approximation’) ignores all correlations between model parameters, resulting in a significant underestimation of uncertainties in the posterior distribution [22, 50]. To avoid these two extremes, PSVI focuses on capturing the most important (dominant) posterior correlations in the model vector $\mathbf{m}$, guided by prior knowledge of the physics controlling typical imaging inverse problems. Specifically, our implementation of PSVI includes posterior correlations between pairs of locations that are in spatial proximity to each other, typically within one dominant wavelength in FWI problems, since wavefield reflection and refraction respond to physical properties of the subsurface averaged spatially over $\sim$1 wavelength. All other parameter correlations are ignored. This structure of posterior correlations has been observed to emerge naturally in a number of previous nonlinear studies, validating this approach [52].
We implement PSVI by defining specific off-diagonal elements of $\mathbf{L}$ to be real number parameters to be optimised during inversion; all other elements are set to be zero, thus assuming spatial independence of the corresponding parameter pairs [52]. This results in a sparse structure for $\mathbf{L}$, and thus also the covariance matrix $\boldsymbol{\Sigma}$, requiring a manageable number of hyperparameters to be optimised in the inversion while still capturing what are typically found to be the most significant correlations.
We acknowledge that this approach does ignore some inter-parameter correlations which are not strictly zero. However, now that we have managed to perform fully nonlinear Bayesian FWI in previous work [e.g., 52], we observe clearly that the dominant posterior correlations almost always occur locally, between cells that are up to a wavelength or so apart both horizontally and vertically. This is because within a wavelength there is little ability to resolve much more than the average velocity values, so sub-wavelength velocities can be varied such that they preserve the correct mean value without compromising data fit. Correlations with parameters in cells that are further afield break down very rapidly, because those parameters are also most strongly correlated with values in cells in their local neighbourhoods. In other words, while it is true that there are far field correlations, they are demonstrably (and intuitively) far less important than those in local neighbourhoods. In 3D FWI problems we are obliged to make some approximation to reduce memory requirements when constructing a covariance matrix. To illustrate, in the following section we present a 3D FWI example with a grid size of $101\times 101\times 63$. If we entertain even only single-precision floating-point format (each element has a size of 4 bytes), storing a full covariance matrix requires about 1 TB memory, which is infeasible for most cases in reality.
@@
Considering the computational cost of 3D Bayesian FWI, a top priority is to improve its efficiency. PSVI does so at the cost of some loss of generality in the solution found due to the Gaussian foundations for the posterior distribution. Zhao and Curtis [52] demonstrated that the method nevertheless produces reasonable statistical information about the full, nonlinear posterior distribution. We therefore accept the loss of generality and employ the method in this work.
@@ 2.3 Variational Prior Replacement
The prior replacement was developed for situations in which we wish to calculate various Bayesian solutions to an inverse problem using different classes of prior information given the same observed dataset [42]. The variational version of the method used here, called variational prior replacement (VPR), was developed to improve its computational efficiency [53]. Say we have two different prior pdf’s $p_{old}(\mathbf{m})$ and $p_{new}(\mathbf{m})$ (hereafter subscripts old and new are used to denote the order in which the pdf’s are estimated or used). According to Bayes’ rule (equation 1), the two posterior probability distributions can be calculated by:
@@
In equation 4, the (old) Bayesian problem given data $\mathbf{d}_{obs}$ and prior pdf $p_{old}(\mathbf{m})$ is solved using variational inference, and the solution $p_{old}(\mathbf{m}|\mathbf{d}_{obs})$ is approximated by the variational distribution $q_{old}(\mathbf{m})$. The third and forth lines in equation 5 are obtained by substituting equation 4 into the second line. Term $p_{new}(\mathbf{m})/p_{old}(\mathbf{m})$ updates the solution that includes the old prior to a solution that contains the new one, and we set constant $k=p_{old}(\mathbf{d}_{obs})/p_{new}(\mathbf{d}_{obs})$. This implies that the new posterior distribution can be calculated from the old one by updating prior information post inversion, as shown by Walker and Curtis [42]: the inversion corresponding to the new prior distribution does not need to be solved from scratch.
Evaluating $k$ requires the two evidence terms to be calculated which is computationally intractable. Zhao and Curtis [53] introduced a second variational distribution $q_{new}(\mathbf{m})$ to approximate $p_{new}(\mathbf{m}|\mathbf{d}_{obs})$ (fifth line in equation 5) by minimising the KL divergence between the new posterior and $q_{new}(\mathbf{m})$, without requiring the value of $k$ to be known explicitly. This step is called variational prior replacement.
A variety of variational inference methods can be used to calculate $q_{old}(\mathbf{m})$ and $q_{new}(\mathbf{m})$ [22, 46, 56]. In this paper, we employ PSVI because it is efficient, and importantly because its probability value is easy to evaluate for any model parameter values post inversion. Using PSVI to find $q_{old}(\mathbf{m})$ thus allows us to perform VPR efficiently to find $q_{new}(\mathbf{m})$ in the fifth line in equation 5 that approximates the solution for any of the alternative prior distributions.
@@
We consider a synthetic 3D acoustic FWI problem. As displayed in Figure 1a, the true velocity model used in this test is a part of the 3D overthrust model [2], which contains $63\times 101\times 101$ grid cells in Z (depth), Y and X directions, respectively, with a cell size of 50 m in each direction. Figures 1c – 1e show three 2D slices of the true velocity structure at locations Y = 2.5 km, X = 2.5 km, and Z = 1.25 km (horizontal slice), respectively. We deploy 81 impulsive sources (red stars) and 10,201 receivers (white dots) at the surface with spacings of 500 m and 50 m, respectively. Both observed and synthetic data are generated by solving the 3D acoustic wave equation using a time-domain pseudo-analytical method [13] with an Ormsby wavelet [32].
We define a uniform prior distribution for the velocity values at different depths, with upper and lower bounds shown in Figure 1b. We use this uniform prior pdf and a diagonal Gaussian likelihood function (implying uncorrelated data noise - not a requirement of the method) to perform variational Bayesian inversion.
Figure 1: (a) True velocity model. Source and receiver locations are indicated by red stars and white dots at the surface (the latter are so close together that they may be observed as a pale haze). (b) Upper and lower bounds of a uniform prior distribution over seismic velocity at each depth. (c), (d) and (e) Two vertical and one horizontal 2D slices of the true velocity model at Y = 2.5 km, X = 2.5 km, and Z = 1.25 km, respectively. These three sections are marked by dashed black lines in panel (a). Dashed red lines in (c) display locations of two vertical profiles along which posterior marginals are compared in the main text.
3.1 Comparison with Previous Results
We apply PSVI to the above 3D FWI problem, and compare the inversion results with those obtained from three other methods presented in Zhang et al. [50]: mean field automatic differentiation variational inference [mean field ADVI – 22], Stein variational gradient descent [SVGD – 25], and stochastic SVGD [sSVGD – 16]. The inversion settings in this example are exactly the same as those used in Zhang et al. [50], so we are able to compare the four sets of results directly. We set the initial value of PSVI to be a standard Gaussian distribution as suggested by Kucukelbir et al. [22]. This results in a laterally homogeneous initial mean velocity model, with velocity values at different depths being the mean of the uniform prior distribution in Figure 1b.
Figure 2 displays the inversion results of the vertical slice shown in Figure 1c, obtained using different methods indicated in the title of each panel. From top to bottom, each row shows the posterior mean velocity, standard deviation, and the relative error maps, respectively, where the relative error is the absolute difference between the true and mean velocities divided by the standard deviation at each point. Each mean velocity map recovers the main features of the true velocity model. However, we also observe that some structures are inverted incorrectly, such as those inside the dashed black boxes in Figure 2, which are supposed to be horizontal according to the true velocity model displayed in Figure 1c. This is possibly because all four inversions fail to recover correct low-wavenumber components of the true model due to $2\pi$ phase shifts in forward-modelled waveform data fitting the observed data equally well – a phenomenon often referred to as cycle skipping in FWI. Mean field ADVI provides the lowest posterior uncertainties due to its theoretical assumption of an uncorrelated Gaussian posterior pdf. Uncertainty values from SVGD and PSVI are larger than those from mean field ADVI, and are smaller than those from sSVGD. Most of the relative errors in the four plots in the bottom row remain relatively high since they all underestimate the posterior uncertainties to some extent for such a high dimensional inverse problem.